OpenStudy (anonymous):

Solve the following equation analytically. 2 log7(x) = log7(2) + log7(x + 4)

OpenStudy (anonymous):

from given eqn u can write it like bellow.. x^2=2(x+4) => x^2-2x-4=0

OpenStudy (anonymous):

it will be \[x^2-2x-8\]

OpenStudy (anonymous):